OpenStudy (anonymous):

Two trains leave stations 240 miles apart at the same time and travel toward each other. One train travels at 65 miles per hour while the other travels at 85 miles per hour. How long will it take for the two trains to meet?

OpenStudy (anonymous):

1.6?

Parth (parthkohli):

One train would cover x miles. The other one would cover 240 - x. Time taken for the first train to cover x miles = {x/65} hours Time for the second train to cover 240 - x miles = (240-x)/85 hours Equating both times \[x/65 = (240 - x)/85\] =>  30x = 3120 x = 104 miles So, answer is 1.6 hours